Write down wavefunction for state n=2
l=0 m=0
1
Expert's answer
2020-11-02T10:52:50-0500

Solution

Wavefunction for quantum number n=2 l=0 m=0

"\psi_{200}=\sqrt\frac{z^3}{32\pi a_0^3} (1-\frac{r}{2a_0}) e^{-\frac{r}{2a_0}}"